Status Report

NASA Space Station On-Orbit Status 9 July 2004

By SpaceRef Editor
July 9, 2004
Filed under , , ,
NASA Space Station On-Orbit Status 9 July 2004

SpaceRef note: This NASA Headquarters internal status report, as presented here, contains additional, original material produced by (copyright © 2004) to enhance access to related status reports and NASA activities.

All ISS systems continue to function nominally except those noted previously or below.

Before breakfast and exercise, the crew completed the blood analysis part of the current IMG PHS (Integrated Medical Group/Periodic Health Status) assessment, taking turns at assisting as CMO (Crew Medical Officer) and then at being the examined subject, using the U.S. PCBA (portable clinical blood analyzer).  Clinical evaluation of the PHS protocol followed later.  Third part of the PHS, subjective evaluation by each crewmember, was performed later in the day.  Afterwards, Mike Fincke completed IFEP data entry for both of them and stowed the hardware.   [The PHS exam, with PCBA analysis and clinical evaluation, is guided by special software (IFEP, in-flight examination program) on the medical equipment computer (MEC).  While PCBA analyzes total blood composition, the blood’s red blood cell content (hematocrit) is measured by the complementary Russian MO-10 protocol.]

Fincke and Padalka, in turn, completed the mandatory 30-min. CBT (computer-based training with video and audio) to refresh their CMO proficiency/rating.   [To maintain proficiency in using HMS (health maintenance systems) hardware including ACLS (advanced cardiac life support) in contingency situations where crew life is at risk, these training sessions are performed once a month to review equipment and procedures via CBT.  Besides ACLS, procedures include airway obstruction management, i.e., review of suction device, nasal airway, intubating laryngeal mask airway (ILMA) with endotracheal tube, and cricothyrotomy (incision to re-enable breathing air inflow).]

Previous Reports

ISS On-orbit Status [HQ]
ISS Status [JSC]
Shuttle Processing [KSC]

The crew also completed their first 30-min. fire drill/OBT (on-board training), a mandatory periodic requirement specifically written for the current two-person crew.  Today’s training consisted of a tabletop review of a specific Lab fire case, supported by tagup with emergency response specialists at MCC-H via S-band.   [The fire case reviewed today was a modification and continuation of the false fire event that occurred early in the morning on 5/18/04.]

The Science Officer terminated the recharge of the MedOps defibrillator equipment’s first battery, begun yesterday, and then started the procedure on battery #2.  Later in the day, he terminated the charging and measured the battery voltage with the Scopemeter.

Mike also floated into the U.S. Airlock (A/L) for the periodic inspection of its single lighting unit’s ELPS (emergency lighting power supply).

Foreshadowing the trash packing for Progress 14P, Gennady tagged up with IMS (inventory management system) specialists at TsUP/Moscow to discuss his upcoming inventory audit of Russian health maintenance system items (SGO), used medical support items (clothes, underwear, countermeasure gear) and sanitary-hygiene equipment (SLG) with expired shelf life.   [In Stage 1 of the inventory, Padalka will open stowage panels, check clothes and hygiene items against an uplinked list and stow found clothes and SLGs in bags for disposal.  Stage 2, the remaining portion of the clean-up work, will be performed in September.]

In preparation for Fluid Merging Viscosity Measurement (FMVM) payload activities on Monday (7/12), Mike discussed the experiment with its Principal Investigator (PI) and team members in a 15-min. teleconference.  He then set up the MWA (Maintenance Work Area) containment system in the Lab, a portable workbench with a tabletop that measures 36 by 25 inches.   [Understanding the viscosity of molten materials is important for everything from designing laboratory experiments to industrial production of materials.  One way to determine viscosity is to measure how long it takes two spheres of liquid to merge into a single spherical drop: on contact a neck will from between the two drops, increasing in thickness until the two drops become one single sphere.  On Earth, gravity distorts liquid spheres, and drops are too heavy to be supported by strings.  Drop distortion should not occur in the ISS’s microgravity environment, and the drops can be held on strings.  To verify this technique as an accurate viscosity measurement method, the FMVM experiment uses fluids with known viscosities: honey, corn syrup, glycerin and silicone oil.  Several runs will be conducted — some with equal diameter drops and others with different size drops.  The crewmember releases two drops from a syringe onto strings and records digital images of the drops as they coalesce to form one drop.  The initial diameters of the drops will be measured.]

Padalka tagged up with the PI and Co-Investigator of the ADUM (Advanced Diagnostic Ultrasound in Micro-G) experiment for an analysis and discussion of yesterday’s “Scan B” abdominal scanning session on Mike Fincke and the “Scan Z” bone scans taken the day before.

The crew conducted the periodic inspection of station bacterial filters and smoke detectors in the A/L, Lab and Node.   [There are one bacterial filter and smoke detector (SD) in the A/L, plus two SDs and one filter each in the Node and Lab.]

Gennady conducted the periodic (monthly) functional closure test of the Vozdukh CO2 removal system’s emergency vacuum valves (AVK, last time done: 6/28).   [The AVKs are critical because they close the Vozdukh’s vacuum access lines in the event of a malfunction in the regular vacuum valves (BVK) or a depressurization in the Vozdukh valve panel (BOA).  Access to vacuum is required to vent carbon dioxide (CO2) during the regeneration of the absorbent cartridges (PP).  During nominal operation, the AVK valves remain open.]

The FE/SO performed the regular once-a-week maintenance reboot on the operational PCS (portable computer system) laptops and the restart of the OCA comm router laptop (every two weeks).  [The PCS in the Service Module (SM) suffered a hard drive failure early yesterday, and for ~15 min. the station had only one PCS laptop operational, until the crew had swapped out the failed drive with a spare to restore Flight Rule requirement.]

Fincke activated the EXPRESS rack 2 (ER2) laptop after connecting it with a cable obtained in the morning from ER3.  ER2 will be active overnight and powered down again tomorrow.

Gennady completed the routine maintenance of the SM’s SOZh environment control & life support system and prepared the regular IMS “delta” file for IMS database update, while Mike conducted the regular routine status checkup of autonomous Lab payloads.

Padalka also completed the periodic replenishing of the Elektron’s water supply for electrolysis, filling the KOV thermal loops’ EDV container with purified (deionized) water from the BKO multifiltration/purification column unit.  (Last time done: 7/2)   [The procedure is specifically designed to prevent air bubbles from getting into the BZh liquid unit where they could cause micropump impeller cavitation and Elektron shutdown, as in the past on numerous times.  In the procedure, the EDV water is drawn from the BKO and the air/liquid separator unit (GZhS) while the crewmember checks for any air bubbles in the EDV (and, if visible, estimates their number).]

Mike Fincke had his weekly PFC (private family conference) via Ku-band/video and S-band/audio.

At about 3:10pm EDT, the crew will conduct their weekly teleconference with the ISS Flight Director at MCC-H.

Gennady and Mike completed their full daily physical exercise program on TVIS treadmill, RED resistive expander and VELO bike with force loader.

Yesterday morning, smoke detector #10 in the FGB went off twice, with warning light and tone triggered.  U.S. segment auto isolation response was executed.  Crew and telemetry confirmed that both events were false alarms, and all systems were restored by the ground.

Starting at 5:00am EDT this morning and running for seven hours, MCC-H and its support group in Moscow (HSG) conducted another BCC (backup control center) dry run in test mode, with no involvement of the ISS crew or vehicle.  Purpose of the periodic exercise is to demonstrate BCC functionality and provide proficiency training for HSG (Houston Support Group) personnel at the TsUP/Moscow HSR (Houston Support Room).

Update on EMU troubleshooting:  The testing on 7/8 confirmed that the pumps in both spacesuits (3013 & 3005) are seized.  There is verbal agreement from the Russian side to fly EMU hardware on next month’s Progress 15P mission.  An R&R of EMU pumps or pump assemblies has never been attempted in space.  Ground specialists are drafting a plan for further testing required to understand the rest of the “fault tree legs”.

Update on Lab science window:  A detailed ULD (ultrasound leak detection) inspection of the Lab nadir window has now been added to the crew’s “job jar” task list.  The inspection will verify and further characterize minute leak points discovered around the window frame on 4/23.   [The procedure will verify old leak points and mark new leaks (if any).  Additionally, the task will determine which interface these leaks are across, i.e., either from the redundant pressure pane frame to the hull (leak out to space) or from the primary pressure pane frame to the redundant pressure pane frame (leak into “Volume D” interpane space.)]

Today’s CEO (crew earth observations) photo targets, in the current LVLH attitude no longer limited by flight rule constraints on the use of the Lab nadir/science window, except for the shutter closure and condensation-prevention plan (limited to 90 min. in 24 hours), were Internal waves, Patagonian Shelf (this may have been the crew’s best pass of several today over this target area.  Clouds were expected to remain 50-60 miles offshore with good glint over the continental shelf from San Jorge Gulf to Blanca Bay), and Internal waves, Tuamotu Archipelago (this was the best pass of two over this target region today. Looking left of track for glint and internal waves among the islands and atolls of the Austral Tuamotu Archipelago.)

CEO images can be viewed at these websites:

See also the website “Space Station Challenge” at:

To view the latest photos taken by the expedition 9 crew visit:

Previous NASA ISS On-orbit Status Reports can be found here. Previous NASA Space Station Status Reports can be found here. Previous NASA Space Shuttle Processing Status Reports can be found here. A collection of all of these reports and other materials relating to Return to Flight for the Space Shuttle fleet can be found here.

U.S. & Russian Segment Status  (as of today, 12:56pm EDT)

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is On (16A, =lowest setting).  Vozdukh CO2 scrubber is On.  U.S. CDRA CO2 scrubber is Off.  TCCS (trace contaminant control subsystem) is operating.  SM Gas Analyzer has been calibrated and is used for ppO2 & ppCO2 monitoring.  MCA (major constituents analyzer) is in Life Extending Mode (LEM).  BMP Harmful Impurities unit: absorbent bed #1 in Purify mode, bed #2 in Purify mode.  RS air conditioner SKV-1 is Off, SKV-2 is Off (SM panel mods completed 4/8; SKV-2 activation failed 4/20; is still considered failed).  SFOG slot #2 fan suspect (not usable).
  • SM Working Compartment:  Pressure (mmHg) — n/a; temperature (deg C) — 25.9; ppO2 (mmHg) — 172.7; ppCO2 (mmHg) — 3.7.
  • SM Transfer Compartment:  Pressure (mmHg) — 751; temperature (deg C) — 20.4.
  • FGB Cabin:  Pressure (mmHg) — 756; temperature (deg C) — 21.3.
  • Node:  Pressure (mmHg) — 752.2; temperature (deg C) — 22.5 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • U.S. Lab:  Pressure (mmHg) — 753.7; temperature (deg C) — 24.7; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • Joint Airlock (Equip. Lock):   Pressure (mmHg) — 753.9; temperature (deg C) — 25.9; shell heater temp (deg C) — n/a, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • (n/a = data not available)
  • PMA-1:  Shell heater temp (deg C) — n/a
  • PMA-2:  Shell heater temp (deg C) — n/a.

Electrical Power Systems (EPS):
Both P6 channels fully operational.  BGA (beta gimbal assembly) 2B and 4B both in directed position (Dual angle/”blind” mode, non solar-tracking, biased for drag reduction).
SM batteries:  Battery #6 is in “Cycle” mode; all other batteries (7) are on line in “Partial Charge” mode.  
FGB batteries:  Battery #4 is off line; battery #3 is in “Cycle” mode; all other batteries (4) are on line in “Partial Charge” mode.

Plasma Contactor Unit PCU-1 is in Standby mode; PCU-2 is in Standby mode.

Command & Data Handling Systems:

  • C&C-1 MDM is prime, C&C-2 is backup, and C&C-3 is in standby.
  • GNC-2 MDM is prime; GNC-1 is backup.
  • INT-2 is operating; INT-1 is Off.
  • EXT-1 is On (primary), EXT-2 is Off (backup).
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • PL-1 MDM is Off; PL-2 MDM is Operational.
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 2 redundant lanes (of 3) operational (string 1 dropped out 11/22/03).
  • SM Central Computer (TsVM): 3 redundant lanes (of 3) operational.
  • FGB MDM-1 is Off (failed, 11/21/03); MDM-2 is Operational.

Propulsion System:

  • Total propellant load available: 3928 kg (8660 lb) as of 7/1/04;  [SM(552) + FGB(2772) + Progress M(639)].  (Capabilities: SM — 860 kg; FGB — 6120 kg).

Attitude Control Systems:

  • 3 CMGs on-line (CMG-1 failed, since 6/6/02; CMG-2’s RPC-17 failed 4/21/04; was replaced 6/30/04).
  • State vector source — U.S. SIGI-1 (GPS)
  • Attitude source — U.S. SIGI-1 (GPS)
  • Angular rate source — RGA-1

Flight Attitude:

  • LVLH XVV (local vertical/local horizontal = “earth-fixed”: z-axis in local vertical, x-axis in velocity vector [yaw: -10 deg, pitch: -9 deg, roll: 0 deg]), with CMG/TA (thruster assist) Momentum Management, until 6/28, following the EVA.

Communications & Tracking Systems:

  • FGB MDM-1 is powered Off; FGB MDM-2 is operational.
  • All other Russian communications & tracking systems are nominal.
  • S-band is operating nominally (on string 2).
  • Ku-band is operating nominally (may require a mask).
  • Audio subsystem is operating nominally (IAC-1 is prime, IAC-2 is off).
  • Video subsystem operating nominally.
  • HCOR (high-rate communications outage recorder) is operating nominally.


  • SRMS/Canadarm2 based at Lab PDGF/LEE A, operational on redundant string, off on prime.
  • MBS: KA (keep alive) power on both strings. 
  • MT: latched and mated at WS4. 
  • POA: KA power on both strings.
  • RWS (robotics workstations): Lab RWS is On (DCP connected); Cupola RWS is Off.

ISS Location NOW

Full Size/Update
Real Time ISS TrackerMore Links

ISS Orbit  (as of this morning, 6:48am EDT [= epoch]):

  • Mean altitude — 359.9 km
  • Apogee height — 364.0 km
  • Perigee height — 355.9 km
  • Period — 91.7 min.
  • Inclination (to Equator) —  51.6304 deg
  • Eccentricity — 0.0005977
  • Solar Beta Angle — 69.7 deg
  • Orbits per 24-hr. day — 15.70
  • Mean altitude loss in last 24 hours — 28 m (!)
  • Revolutions since FGB/Zarya launch (Nov. ’98)  — 32181

ISS Altitude History

Apogee height Mean AltitudePerigee height

ISS Altitude History

Additional satellite tracking resources can be found at
For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see In addition, information on International Space Station sighting opportunities can be found at on NASA’s Human Spaceflight website. The current location of the International Space Station can be found at at NASA’s Marshall Space Flight Center.

SpaceRef staff editor.